# Experimental Validation of Reflective Near-Field Beamfocusing using a $b$-bit RIS

### v1

@@ Abstract
This paper presents the first experimental validation of reflective near-field beamfocusing using a reconfigurable intelligent surface (RIS). While beamfocusing has been theoretically established as a key feature of large-aperture RISs, its practical realization has remained unexplored. We derive new analytical expressions for the array gain achieved with a $b$-bit RIS in near-field line-of-sight scenarios, characterizing both the finite depth and angular width of the focal region. The theoretical results are validated through a series of measurements in an indoor office environment at 28 GHz using a one-bit 1024-element RIS. The experiments confirm that near-field beamfocusing can be dynamically achieved and accurately predicted by the proposed analytical model, despite the presence of hardware imperfections and multipath propagation. These findings demonstrate that near-field beamfocusing is a robust and practically viable feature of RIS-assisted wireless communications.
@@ I Introduction
Reconfigurable intelligent surfaces (RISs) have emerged as a novel kind of relay that, through their reflections, can simultaneously amplify signals in desired directions and change the shape of wavefronts [1]. An RIS consists of a two-dimensional array of subwavelength-sized elements with reconfigurable reflection coefficients, which can be tuned to modify how the incident waves are scattered. The previous literature contains plenty of algorithms for selecting the RIS configuration to enhance communication capacity-related metrics by shaping the propagation environment [2]. Several key RIS features have been validated experimentally, including the cascaded pathloss formula [3], over-the-air RIS configuration in outdoor scenarios [4], the angular beampatterns [5, 6], and how an RIS can encode information by switching its configuration [7].
@@
In this paper, we address this gap through a dedicated experimental campaign in a real indoor environment. In particular, we study finite-depth beamfocusing with a $b$-bit RIS by first deriving novel theoretical array gain expressions and then comparing them to measurement results. We demonstrate that beamfocusing is not a theoretical curiosity, but can be utilized in the studied office scenarios, and also that the RIS resolution does not widen the depth of focus.
@@ II System Model and Ideal RIS Configuration
where $s$ is the transmitted signal and $n$ is the receiver noise. The end-to-end channel coefficient $h\in\mathbb{C}$ consists of two components. The static paths that does not involve the RIS are jointly represented by $h_{\textrm{s}}$, while the reconfigurable path via the RIS is $\mathbf{h}_{\textrm{r}}^{\mbox{\tiny$\mathrm{T}$}}\mathbf{\Phi}\mathbf{h}_{\textrm{t}}$, where $\mathbf{h}_{\textrm{t}}=[h_{\textrm{t},1},\ldots,h_{\textrm{t},N}]^{\mbox{\tiny$\mathrm{T}$}}\in\mathbb{C}^{N}$ is the channel from the transmitter to the RIS, $\mathbf{h}_{\textrm{r}}=[h_{\textrm{r},1},\ldots,h_{\textrm{r},N}]^{\mbox{\tiny$\mathrm{T}$}}\in\mathbb{C}^{N}$ is the channel from the RIS to the receiver, and $\bm{\Phi}=\mathrm{diag}\left(e^{j\phi_{1}},\ldots,e^{j\phi_{N}}\right)$ is the diagonal matrix containing the $N$ reconfigurable phase-shifts in the RIS.
@@ III Near-Field Beamfocusing with a $b$-bit RIS
It would be $N^{2}|\alpha_{\textrm{r}}\alpha_{\textrm{t}}|^{2}$ if $\epsilon_{1}=\ldots=\epsilon_{N}=0$. However, these phase deviations are non-zero when using a practical $b$-RIS. The deviations are deterministic for LOS channels and can be computed for a specific RIS setup; however, since there are many parameters, their joint effect appears random-like. Hence, to quantify their impact, we approximate them as independent random variables that are uniformly distributed on $\mathcal{Q}_{b}$, just as in the independent Rayleigh fading scenario considered in [8, 9]. The average channel gain then becomes
@@ III-A Depth of Beamfocusing
We will now characterize the shape of the beam reflected from the RIS, especially in the near-field, where the depth can be finite. Suppose the RIS is configured to focus its reflected signal at a point in the correct direction, but at a different focal distance $F$ as illustrated in Fig. 1. We want to quantify the resulting loss in channel gain.
@@ III-B Width of Beamfocusing
The beamwidth around the focal point can be analyzed similarly to the depth. Suppose the RIS is configured to focus its reflected signal at a point with the correct distance but incorrect angles. Suppose the RIS configuration is selected as in (5) but with $\varphi_{\textrm{r}},\theta_{\textrm{r}}$ replaced by $\varphi_{\textrm{F}},\theta_{\textrm{F}}$, which is a point on the curve in Fig. 1. The channel gain at the receiver then becomes
@@
The dependence on the angle mismatches appears through $B$ and $C$, which are defined in (17)–(18). This expression has a sinc-squared behavior, which is relatively common in beamwidth analysis involving UPAs (cf. [12, Eq. (32)]). Interestingly, when $N$ is large, the bit-resolution does not affect the beamwidth, which is determined by the nulls of the last two sinc-squared functions in (21). Hence, even if the hardware resolution limits the maximum array gain, it does not change the shape of the beam. We will validate this expression experimentally in Section IV.
@@ IV Experimental Validation
The phase configuration $\bm{\Phi}=\mathrm{diag}\left(e^{j\phi_{1}},\ldots,e^{j\phi_{N}}\right)$ of the RIS is determined by the TMYTEK software, which takes angles and distances to the transmitter and receiver (with respect to the RIS) as input. The exact calculation is not disclosed in the user manual, but we believe it aligns the phases similarly to (2) with $h_{\textrm{s}}=0$ and then quantizes the results to the nearest implementable value. The software might also compensate for mutual coupling within the RIS array.
@@ IV-A First Measurement Setup
In Fig. 3, we show the array gain (normalized to have the peak value $1$) as a function of the focal distance $F$ when the receiver distance is $d_{\textrm{r}}=0.6$ m. We compare the measured values with the theoretical formula in (14) and with a simulation where the phase-shifts were computed using (2) and then quantized to the nearest points in $\mathcal{Q}_{1}$. The measured and theoretical curves show good agreement, with both exhibiting a sharp focusing effect around $F=d_{\textrm{r}}$. This demonstrates that the novel formula properly captures the array gain loss caused by phase quantization. The half-power beamdepth is around $1$ m. The dash-dotted black simulation curve fluctuates rapidly when the RIS is defocused, due to the quantization effect in the 1-bit phase-shifts. The solid blue curve, based on the formula in (14), appears at the center of these fluctuations. This validates that this approximate formula accurately represents the average array gain.
@@
In Fig. 4, we increase the receiver distance to $d_{\textrm{r}}=1.2$ m, while keeping all other parameters constant. The theoretical and experimental results maintain consistent focusing characteristics, and the measured peak aligns closely with the theoretical focus. Interestingly, the measured curve exhibits a slightly narrower beamdepth than the theoretical curves. This indicates that practical characteristics not considered in our theoretical model (e.g., mutual coupling, gain variations across elements, and additional propagation paths) will not destroy the beamfocusing ability of the RIS, but rather strengthen it.
@@ IV-B Second Measurement Setup
Fig. 5 shows how the array gain varies when the focal point is moved in both distance and azimuth angle, as compared to the boresight receiver at $d_{\textrm{r}}=0.8$ m. This corresponds to moving along the two axes shown in Fig. 1. All the intersection points in this graph represent the measurements, while the surface indicates the shape. The yellow peak value coincides with the true location of the receiver, and there is a roughly ellipsoidal area around it with a substantial array gain (shown in yellow and green). This result demonstrates that one can achieve beamfocusing with a textbook-like shape, as sketched in Fig. 1, in a real office environment.
Next, we take a closer look at the beamwidth behavior by varying the azimuth angle of the focal point in the range $20^{\circ},20^{\circ}]$, while keeping the focal distance and elevation angle the same as for the receiver. Fig. 6 shows the measurement results and two theoretical curves. The solid blue curve is computed using the formula in (21) and matches the measurements closely, particularly within the mainlobe. The half-power beamwidth is around $3$ degrees. The black dash-dotted curve is once again computed by calculating theoretical LOS channels and quantizing the ideal RIS phase-shifts. This simulation curve is close to the formula-based curve, but when it deviates in the sidelobes, it is slightly closer to the measurements.
@@
There is a minor discrepancy between measurements and theory around $15^{\circ}$. In this case, the RIS reflects the signal back toward the transmitter, and the metal object that we used to block the direct path between the transmitter and receiver is likely reflecting this signal toward the receiver.
@@ V Conclusions
The near-field communication literature frequently highlights finite-depth beamfocusing as an attractive new interference-suppression feature that concentrates radiated signals toward a spatial region confined in both angle and depth. In this paper, we have demonstrated that this theoretically appealing concept can be realized in practical indoor environments using a dynamic $b$-bit RIS. We derived novel analytical expressions for the depth and width of the near-field reflective beamfocusing effect and validated them experimentally in 28 GHz measurement setups, considering two different office environments. The close agreement between theoretical predictions and measured results confirms that the derived models accurately capture real-world behavior, despite neglecting mutual coupling, element gain variations, and multipath propagation. To the best of our knowledge, this paper provides the first experimental validation of reflective near-field beamfocusing with a dynamic RIS, establishing its feasibility for future RIS-assisted communication systems.

### v2

@@ Abstract
This paper presents the first experimental validation of reflective near-field beamfocusing using a reconfigurable intelligent surface (RIS). While beamfocusing has been theoretically established as a key feature of large-aperture RISs, its practical realization has remained unexplored. We derive new analytical expressions for the array gain achieved with a $b$-bit RIS in near-field line-of-sight scenarios, characterizing both the finite depth and angular width of the focal region. The theoretical results are validated through a series of measurements in an indoor office environment at 28 GHz using a one-bit 1024-element RIS. The experiments confirm that beamfocusing can be dynamically achieved and accurately predicted by the proposed simplified analytical model, despite the presence of hardware imperfections and multipath propagation. These findings demonstrate that near-field beamfocusing is a robust and practically viable feature of RIS-assisted wireless communications.
@@ I Introduction
Reconfigurable intelligent surfaces (RISs) have emerged as a new kind of repeater that through their reflections can simultaneously amplify signals in desired directions and change the shape of wavefronts [1]. An RIS consists of a two-dimensional array of subwavelength-sized elements with reconfigurable reflection coefficients, which can be tuned to modify how the incident waves are scattered. The previous literature contains plenty of algorithms for selecting the RIS configuration to enhance communication capacity-related metrics by shaping the propagation environment [2]. Several key RIS features have been validated experimentally, including the cascaded pathloss formula [3], over-the-air RIS configuration in outdoor scenarios [4], the angular beampatterns [5, 6], and how an RIS can encode information by switching its configuration [7].
@@
In this paper, we address this gap through a dedicated experimental campaign in a real indoor environment. In particular, we study finite-depth beamfocusing with a $b$-bit RIS by first deriving novel theoretical array gain expressions and then comparing them to measurement results. We demonstrate that beamfocusing is not a theoretical curiosity, but can be achieved in the studied office scenarios, that a limited RIS resolution does not widen the depth of focus, and that simple models are sufficient to predict the behavior of practical hardware.
@@ II System Model and Ideal RIS Configuration
where $s$ is the transmitted signal and $n$ is the receiver noise. The end-to-end channel coefficient $h\in\mathbb{C}$ consists of two components. The static paths that does not involve the RIS are jointly represented by $h_{\textrm{s}}$, while the reconfigurable path via the RIS is $\mathbf{h}_{\textrm{r}}^{\mbox{\tiny$\mathrm{T}$}}\mathbf{\Phi}\mathbf{h}_{\textrm{t}}$, where $\mathbf{h}_{\textrm{t}}=[h_{\textrm{t},1},\ldots,h_{\textrm{t},N}]^{\mbox{\tiny$\mathrm{T}$}}\in\mathbb{C}^{N}$ is the channel from the transmitter to the RIS, $\mathbf{h}_{\textrm{r}}=[h_{\textrm{r},1},\ldots,h_{\textrm{r},N}]^{\mbox{\tiny$\mathrm{T}$}}\in\mathbb{C}^{N}$ is the channel from the RIS to the receiver, and $\bm{\Phi}=\mathrm{diag}\left(e^{j\phi_{1}},\ldots,e^{j\phi_{N}}\right)$ is the diagonal matrix containing the $N$ reconfigurable phase-shifts in the RIS. Note that this model neglects mutual coupling and nonlinearities, and we will later show that it is anyway sufficiently accurate when studying practical RIS that are configured to mitigate these effects.
@@ III Near-Field Beamfocusing with a $b$-bit RIS
It would equal $N^{2}|\alpha_{\textrm{r}}\alpha_{\textrm{t}}|^{2}$ if $\epsilon_{1}=\ldots=\epsilon_{N}=0$. However, these phase deviations are non-zero when using a practical $b$-bit RIS. The deviations are deterministic for LOS channels and can be computed for a specific RIS setup; however, since there are many parameters, their joint effect appears random-like. Hence, to quantify their impact, we approximate them as independent random variables that are uniformly distributed in $\mathcal{Q}_{b}$, just as in the independent Rayleigh fading scenario considered in [8, 9]. The average channel gain then becomes
@@ III-A Depth of Beamfocusing
We will now characterize the shape of the beam reflected from the RIS, especially in the radiative near-field, where the depth can be finite. Suppose the RIS is configured to focus its reflected signal on a point in the correct direction, but at a different focal distance $F$ as illustrated in Fig. 1. We want to quantify the resulting loss in channel gain analytically.
@@ III-B Width of Beamfocusing
The beamwidth around the focal point can be analyzed similarly to the depth. Suppose the RIS is configured to focus its reflected signal on a point with the correct distance but incorrect angles. Suppose the RIS configuration is selected as in (5) but with $\varphi_{\textrm{r}},\theta_{\textrm{r}}$ replaced by $\varphi_{\textrm{F}},\theta_{\textrm{F}}$, which is a point on the curve in Fig. 1. The channel gain at the receiver then becomes
@@
The dependence on the angle mismatches appears through $B$ and $C$, which are defined in (17)–(18). This expression has a sinc-squared behavior, which is relatively common in beamwidth analysis involving UPAs (cf. [12, Eq. (32)]). Interestingly, when $N$ is large, the bit-resolution does not affect the beamwidth, which is determined by the nulls of the last two sinc-squared functions in (21). Hence, even if the hardware resolution limits the maximum array gain, it does not change the shape of the beam. We will validate this expression experimentally in the next section.
@@ IV Experimental Validation
The phase configuration $\bm{\Phi}=\mathrm{diag}\left(e^{j\phi_{1}},\ldots,e^{j\phi_{N}}\right)$ of the RIS is determined by the TMYTEK software, which takes angles and distances to the transmitter and receiver (with respect to the RIS) as input. The exact calculation is not disclosed in the user manual, but we believe it aligns the phases similarly to (2) with $h_{\textrm{s}}=0$ and quantizes the results to the nearest implementable value. The software compensates for mutual coupling and other RIS hardware impairments.
@@ IV-A First Measurement Setup
In Fig. 3, we show the array gain (normalized to have the peak value $1$) as a function of the focal distance $F$ when the receiver distance is $d_{\textrm{r}}=0.6$ m. We compare the measured values with the theoretical formula in (14) and with a simulation where the phase-shifts were computed using (2) and then quantized to the nearest points in $\mathcal{Q}_{1}$. The measured and theoretical curves show good agreement, with both exhibiting a sharp focusing effect around $F=d_{\textrm{r}}$. This demonstrates that our new formula properly captures how the array gain depends on the distance under 1-bit quantization. The half-power beamdepth is around $1$ m. The dash-dotted black simulation curve fluctuates rapidly when the RIS is defocused, due to the quantization effect in the 1-bit phase-shifts. The solid blue curve, based on the formula in (14), appears at the center of these fluctuations. This validates that this formula accurately approximates the average array gain.
@@
In Fig. 4, we increase the receiver distance to $d_{\textrm{r}}=1.2$ m, while keeping all other parameters constant. The theoretical and experimental results maintain consistent focusing characteristics, and the measured peak aligns closely with the theoretical focus. Interestingly, the measured curve exhibits a slightly narrower beamdepth than the theoretical curves. This indicates that practical characteristics not considered in our system model (e.g., mutual coupling, gain variations across elements, and additional propagation paths) will not destroy the beamfocusing ability of the RIS, but rather strengthen it.
@@ IV-B Second Measurement Setup
Fig. 5 shows how the array gain varies when the focal point is moved in both distance and azimuth angle, as compared to the boresight receiver at $d_{\textrm{r}}=0.8$ m. This corresponds to moving along the two axes shown in Fig. 1. All the intersection points in this graph represent the measurements, while the surface and color indicate the shape. The yellow peak value coincides with the true location of the receiver, and there is a roughly ellipsoidal area around it with a substantial array gain (shown in yellow and green). This result demonstrates that one can achieve beamfocusing with a textbook-like shape, as sketched in Fig. 1, in a real office environment.
Next, we take a closer look at the beamwidth behavior by varying the azimuth angle of the focal point in the range $20^{\circ},20^{\circ}]$, while keeping the focal distance and elevation angle the same as for the receiver. Fig. 6 shows the measurement results and two theoretical curves. The solid blue curve is computed using the formula in (21) and matches the measurements closely, particularly within the mainlobe. The half-power beamwidth is around $3$ degrees. The black dash-dotted curve is once again computed by calculating theoretical LOS channels and quantizing the ideal RIS phase-shifts. This simulation curve is close to the formula-based curve, but when it deviates in the sidelobes, it is slightly closer to the measurements. There is a minor discrepancy between measurements and theory around $15^{\circ}$. In this case, the RIS reflects the signal back toward the transmitter, and the metal object that we used to block the direct path between the transmitter and receiver reflects this signal toward the receiver.
@@ V Conclusions
The near-field communication literature frequently highlights finite-depth beamfocusing as an attractive new interference-suppression feature that concentrates radiated signals toward a spatial region confined in both angle and depth. In this paper, we have demonstrated that this theoretically appealing concept can be realized in practical indoor environments using a dynamic $b$-bit RIS. We derived novel analytical expressions for the depth and width of the near-field reflective beamfocusing effect and validated them experimentally in 28 GHz measurement setups, considering two different office environments. The close agreement between theoretical predictions and measured results confirms that the derived formulas accurately capture real-world behaviors, even if derived from a system model that neglects mutual coupling, element gain variations, and multipath propagation. Importantly, this indicates that more advanced electromagnetic models are not needed when studying LOS setups where the RIS hardware has been designed to actively compensate for hardware impairments. To the best of our knowledge, this paper provides the first experimental validation of reflective near-field beamfocusing with a dynamic RIS, establishing its feasibility for future RIS-assisted communication systems.